OpenStudy (anonymous):

if cos(t)=0.8 and t is in 3rd quad, find sin(2t) is this 0.96 for the answer?

OpenStudy (anonymous):

2sin(x)cos(x) = sin(2x) so 2sin(x)(0.8) = sin(2x) so you need to find cos(x) so we use the definition cos^(2)(x) + sin^(2)(x) = 1 sub 0.8 into the formula and you should get something like this: \[\cos(x) = \pm \sqrt{1-(0.8)^2}\] Now to remember what is positive and what is negative in the quadrants remember, All Students Take calculus". Because QI = (A)ll positive, QII = (S)ine, csc is positive QIII = (T)an cot is positive QIV = Cos and sec are positive so sin would be negative. Follow these steps and you should be fine